1 Problem 21.44 Problem 21.44 Part A A 93 Ω resistor has a current of 0.65 A and is connected in series with a 140 Ω resistor. | What is the emf of the battery to which the resistors are connected? Express your answer using two significant figures kV Submit My Answers Give Up
Answer #1

I = 0.65 A

equivalent resistance Requ = 93 + 140 = 233 ohm

emf = I R = 233 * 0.65

emf = 151.45 V = 0.151 kV